The probability that A hits a target is 2/5 and the probability that B hits the target is 2/5. Both shoot at the target. Find the probability that at least one of them hits the target, i.e., that A or B (or both) hit the target



1
Expert's answer
2021-12-13T07:28:53-0500

Let A - "first hit"

B - "second hit", then

"P(A\cup B)=1-P(\neg (A\cup B))=1-P(\neg A\cap \neg B)=1-P(\neg A)*P(\neg B)=1-{\frac 3 5}*{\frac 3 5}=1-{\frac 9 {25}}=0.64"